How to solve (2/3)^-2
Answer by Theo(13342)

the easiest way to solve this is to follow the rule that states:

(a/b)^-c = 1 / (a/b)^c

then there is the rule that states:

(a/b)^c = a^c / b^c

there is also the rule that states:

1 / (a^c / b^c) = b^c / a^c

using these three rules, you would get:

(2/3)^2 = 1 / (2/3)^2 = 1 / (2^2 / 3^2) = 3^2 / 2^2 = 9/4 = 2.25

here's a reference on exponent arithmetic rules.

https://www.mathsisfun.com/algebra/exponent-laws.html

notice, he did not give you (x/y)^-n.

however, he did give you x^-n = 1/x^n

the x in this case, doesn't just refer to a variable.
it also refers to an expression.

in this case x represents the expression of (a/b).

when he says x^n = 1/x^-n, when you replace x with a/b, you get (a/b)^-n = 1/(a/b)^n

the a/b within parentheses makes that expression act like a variable until you break it out of the parentheses.

(a/b)^-n = 1/(a/b)^n = 1/(a^n/b^n) = b^n / a^n

here's another reference that is more detailed and also more informative.

https://www.purplemath.com/modules/exponent.htm